# Matrix systems, algebras, and open maps

### v1

@@ 9. Continuity in quantum information theory
The commutative diagram shows that the fiber $\pi_{\mathfrak{g}}^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$ is the set of all $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ such that $\operatorname{red}_{\mathfrak{g}}(\rho)=\operatorname{red}_{\mathfrak{g}}(\sigma)$. Thm. 9.2 proves that the pullback $\Xi_{\tau}:=\Psi_{\tau}\circ\operatorname{red}_{\mathfrak{g}}^{-1}$ of the inference map $\Psi_{\tau}$ under $\operatorname{red}_{\mathfrak{g}}^{-1}$ is continuous at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ if and only if $\operatorname{red}_{\mathfrak{g}}$ is open at $\Xi_{\tau}[(\rho_{\nu})_{\nu\in\mathfrak{g}}]\in\mathcal{D}(\mathcal{A}_{\Omega})$.
A discontinuity of $\Xi_{\tau}$ (with uniform prior $\tau$) was reported for $N=3$ qubits and $\mathfrak{g}:=\{\{1,2\},\{2,3\},\{3,1\}\}$ at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$, where
$\rho_{\nu}:=\frac{1}{2}(\ket{00}\!\!\bra{00}+\ket{11}\!\!\bra{11})\,,\quad\nu\in\mathfrak{g}\,.$
This result by Chen et al. [14, Ex. 4] has an interesting interpretation. The map $\operatorname{red}_{\mathfrak{g}}$ being open [^5] The openness of $\operatorname{red}_{\mathfrak{g}}$ at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a priori weaker than the continuity of $\Xi_{\tau}$ at $\operatorname{red}_{\mathfrak{g}}(\rho)$. The continuity means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ can be matched by an arbitrarily small change of $\rho$ inside the range of $\Xi_{\tau}$ and not just anywhere within $\mathcal{D}(\mathcal{A}_{\Omega})$. Somewhat surprisingly, the two propositions are equivalent by Thm. 9.2. at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ within $\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ is matched by an arbitrarily small change of $\rho$ within $\mathcal{D}(\mathcal{A}_{\Omega})$. Conversely, if the openness fails, then there are arbitrarily small changes of $\operatorname{red}_{\mathfrak{g}}(\rho)$ that are only matched by changes of $\rho$ beyond some positive threshold (in the metric sense). Loosely speaking, a small change of a subsystem abruptly changes the entire system. Such behavior is associated with phase transitions [14]. It motivates every attempt to study the openness problem of $\pi_{\mathfrak{g}}$. This should be done by tuning the interaction pattern $\mathfrak{g}$ to a concrete system. Whether enclosing $\mathcal{R}_{\mathfrak{g}}$ into a *-subalgebra of $\mathcal{A}_{\Omega}$ could simplify the problem, as suggested in Rem. 8.2, is not yet clarified.
Focusing on the continuity of correlations measures, we begin with theory.
@@
Let $\tau:=\mathds{1}_{n}/n$ and let $\mathcal{E}(\mathcal{R}):=\mathcal{E}_{\tau}(\mathcal{R})$ be an exponential family, see Ex. 9.3. The entropy distance from $\mathcal{E}(\mathcal{R})$ is
$d:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho_{1}\mapsto\inf_{\rho_{2}\in\mathcal{E}(\mathcal{R})}S(\rho_{1},\rho_{2})\,.$ (9.6)
Using the von Neumann entropy $S=\phi_{\tau}+\log(n)$, we define
Clearly, the map $\Psi$ equals the inference map $\Psi_{\tau}$ defined earlier in Ex. 9.1. The entropy distance can be written as the difference
@@
between the value of the von Neumann entropy at $\rho$ and its maximal value on the fiber of $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ that contains $\rho$, see p. 1288 in [52].
Formula (9.7) suggests studying the rI-projection [19]
$\Pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\Pi=\Psi\circ\pi\,.$
By (9.3), the range of $\Pi$ is the rI-closure of $\mathcal{E}(\mathcal{R})$. We also consider the restriction $d|_{\mathcal{D}(\mathcal{A})}$ of $d$ and the following restriction of $\Pi$.
$\Pi|_{\mathcal{D}(\mathcal{A})}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{A})\,,\quad\rho\mapsto\Pi(\rho)$
Here, $\mathcal{A}$ is real *-subalgebra of $\mathrm{M}_{n}$ that includes $\mathcal{R}$. The map $\Pi|_{\mathcal{D}(\mathcal{A})}$ is well defined because $\Psi(\mathcal{D}(\mathcal{R}))$ is included in $\mathcal{D}(\mathcal{A})$, as observed in Ex. 9.3 a).
Example 9.6 (Entropy distance II).
The exponential family $\mathcal{E}(\mathcal{R}_{\mathfrak{g}})$ of local Hamiltonians (Ex. 9.4) is interesting because the entropy distance from it quantifies many-body correlations. Amari [4] and Ay [6] studied this type of correlation measure in probability theory. Linden et al. [33] introduced it as a difference of von Neumann entropies like formula (9.7), see also [58, Sec. 1.4.2]. And Zhou [59] proved the equivalence of the two representations (9.6) and (9.7) for density matrices of maximal rank $n$, see also [38]; the equivalence is true without rank restrictions as presented in Ex. 9.5.
We quote from [52, Lemma 5.15 and 4.5] under the assumptions of Ex. 9.5.
@@
Le $\mathcal{A}$ be real *-subalgebra of $\mathrm{M}_{n}$ that includes $\mathcal{R}$.
@@
    For every $\rho\in\mathcal{D}(\mathcal{A})$, the map $\Pi|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\rho$ if and only if $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\rho$.
@@
With reference to Ex. 9.5, we simplify the continuity problem of the entropy distance (and of other invariant maps) using *-subalgebras.
Remark 9.8 (Continuity of the rI-projection).
Let $\mathcal{A}$ be a *-subalgebra of $\mathrm{M}_{n}$ such that $\mathcal{R}\subset\mathcal{A}$. As in Rem. 8.2, the map $\pi$ factors into $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$. Then, for every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the rI-projection $\Pi$ is continuous at $\rho$ if and only if $\Pi|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$.
This assertions needs a justification. By its definition, the map $\Pi=\Psi\circ\pi$ is invariant under $\pi_{\mathcal{A}}$ in the sense that $\Pi=\Pi\circ\pi_{\mathcal{A}}$ holds. This shows $\Pi=\Pi|_{\mathcal{D}(\mathcal{A})}\circ\pi_{\mathcal{A}}$. The claim then follows from Lemma 8.1 b), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\Pi|_{\mathcal{D}(\mathcal{A})}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$. The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm. 1.2.
@@
Rem. 9.8 helps solve the continuity problem of the entropy distance from $\mathcal{E}(\mathcal{R})$ for $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$, based on a solution for $d|_{\mathcal{A}(\mathcal{D})}$ in the algebra $\mathcal{A}:=\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ that was obtained using asymptotic curvature estimates [52].
Real *-subalgebras are excluded from Rem. 9.8 but the conclusion is still true, as Ex. 7.2 can replace Thm. 1.2 in this remark.
@@
and continuous at every point in the complement. (Unitary similarity with respect to $U\oplus 1$ matches $\mathcal{R}$ with the problem considered in [52], as the similarity with respect to $U:=\exp(\mathrm{i}\frac{\pi}{3\sqrt{3}}(X+Y+Z))$ permutes the Pauli matrices $X,Y,Z$ cyclicly.) Lemma 9.7 then shows that the restricted rI-projection $\Pi|_{\mathcal{D}(\mathcal{A})}$ is discontinuous at every point in $\mathcal{G}_{d}$ and continuous at every point in the complement. Now Rem. 9.8 proves that the rI-projection $\Pi$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement. A second application of Lemma 9.7 finally shows that the entropy distance $d$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement.
The points in the set $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ are explicitly described in Ex. 8.3 b). The result is consistent with the inference $\Psi$ being discontinuous at $M$ and continuous at every point different from $M$ (see Ex. 9.3 c)) as per Lemma 9.7 b).
Acknowledgments. The author is grateful to several people for discussions related to this work, among them are Achim Clausing, Chi-Kwong Li, Tim Netzer, Ivan Todorov, and Ilya M. Spitkovsky. The author gave a lecture on the results of this paper in the special session “Numerical Ranges” of the conference IWOTA 2024 in Canterbury, UK, and wishes to thank the organizers for this opportunity.

### v2

@@ 9. Continuity in quantum information theory
The fiber $\pi_{\mathfrak{g}}^{-1}(\sigma)$ over $\sigma\in\mathcal{D}(\mathcal{R}_{\mathfrak{g}})$ is the set of all $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ such that $\operatorname{red}_{\mathfrak{g}}(\rho)=\operatorname{red}_{\mathfrak{g}}(\sigma)$. Thm. 9.2 proves that the pullback $\Xi:=\Psi\circ\operatorname{red}_{\mathfrak{g}}^{-1}$ of the inference map $\Psi$ under $\operatorname{red}_{\mathfrak{g}}^{-1}$ is continuous at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ if and only if $\operatorname{red}_{\mathfrak{g}}$ is open at $\Xi[(\rho_{\nu})_{\nu\in\mathfrak{g}}]\in\mathcal{D}(\mathcal{A}_{\Omega})$.
Chen et al. [13, Ex. 4] discovered a discontinuity of $\Xi$ for $N=3$ qubits and $\mathfrak{g}:=\{\{1,2\},\{2,3\},\{3,1\}\}$ at $(\rho_{\nu})_{\nu\in\mathfrak{g}}\in\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$, where
$\rho_{\nu}:=\frac{1}{2}(\ket{00}\!\!\bra{00}+\ket{11}\!\!\bra{11})\,,\quad\nu\in\mathfrak{g}$
and they offered an interesting interpretation in terms of phase transitions. The map $\operatorname{red}_{\mathfrak{g}}$ being open [^5] The openness of $\operatorname{red}_{\mathfrak{g}}$ at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ is a priori weaker than the continuity of $\Xi$ at $\operatorname{red}_{\mathfrak{g}}(\rho)$. The continuity means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ can be matched by an arbitrarily small change of $\rho$ inside the image of $\Xi$ and not just anywhere in $\mathcal{D}(\mathcal{A}_{\Omega})$. Somewhat surprisingly, the two propositions are equivalent by Thm. 9.2. at $\rho\in\mathcal{D}(\mathcal{A}_{\Omega})$ means that any sufficiently small change of $\operatorname{red}_{\mathfrak{g}}(\rho)$ in $\operatorname{red}_{\mathfrak{g}}[\mathcal{D}(\mathcal{A}_{\Omega})]$ is matched by an arbitrarily small change of $\rho$ within $\mathcal{D}(\mathcal{A}_{\Omega})$. Conversely, if the openness fails, then there are arbitrarily small changes of $\operatorname{red}_{\mathfrak{g}}(\rho)$ that can only be matched by changes of $\rho$ beyond some strictly positive threshold (in the metric sense). Loosely speaking, a small change of a subsystem abruptly changes the entire system. Such behavior is associated with phase transitions. It motivates every attempt to study the openness of $\pi_{\mathfrak{g}}$. This should be done by tuning the interaction pattern $\mathfrak{g}$ to a concrete system. Whether enclosing $\mathcal{R}_{\mathfrak{g}}$ into a *-subalgebra of $\mathcal{A}_{\Omega}$ could simplify this problem, as suggested by Exa. 9.3, is not yet clarified.
We finish this paper with a map whose continuity is more subtle than that of the inference.
@@
The entropy distance from the exponential family $\mathcal{E}(\mathcal{R})$ is defined by
$d:\mathcal{D}(\mathrm{M}_{n})\to\mathbb{R}\,,\quad\rho_{1}\mapsto\inf_{\rho_{2}\in\mathcal{E}(\mathcal{R})}S(\rho_{1},\rho_{2})$ (9.3)
and equals the difference
@@
between the value of the von Neumann entropy at $\rho$ and the maximal value on the fiber of $\pi$ that contains $\rho$, see p. 1288 in [53]. Formula (9.4) suggests studying the continuity of $d$ through the rI-projection
$\Pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathrm{M}_{n})\,,\quad\Pi:=\Psi\circ\pi\,.$
As per Def. 5.2 and equation (D8) in [53], the density matrix $\Pi(\rho_{1})$ is the generalized rI-projection of $\rho_{1}\in\mathcal{D}(\mathrm{M}_{n})$ to $\mathcal{E}(\mathcal{R})$, which is a well-known concept in probability theory [18], and which is defined as follows. A sequence $(\tau_{i})\subset\mathcal{D}(\mathrm{M}_{n})$ rI-converges to $\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$ if $\lim_{i}S(\rho_{2},\tau_{i})=0$ holds. If every sequence $(\tau_{i})\subset\mathcal{E}(\mathcal{R})$ satisfying $\lim_{i}S(\rho_{1},\tau_{i})=d(\rho_{1})$ rI-converges, independently of the sequence, to a unique $\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$, not necessarily in $\mathcal{E}(\mathcal{R})$, then $\rho_{2}$ is the generalized rI-projection of $\rho_{1}$ to $\mathcal{E}(\mathcal{R})$.
We quote from Lemma 5.15 and Lemma 4.5 in [53].
@@
    For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$ the rI-projection $\Pi$ is continuous at $\rho$ if and only if the entropy distance $d$ is continuous at $\rho$.
@@
Example 9.7 (Entropy distance II).
The entropy distance from the exponential family $\mathcal{E}(\mathcal{R}_{\mathfrak{g}})$ of local Hamiltonians (Ex. 9.4) is interesting because it quantifies many-body correlations. Amari [4] and Ay [6] studied this type of correlation measures in probability theory. Linden et al. [33] introduced it to quantum mechanics as a difference of von Neumann entropies like formula (9.4), see also [59, Sec. 1.4.2]. Zhou [60] proved the equality of the two representations (9.3) and (9.4) for density matrices of maximal rank $n$, see also [38]; the equality is true without rank restrictions as stated in Ex. 9.5.
Proposition 9.8.
For every $\rho\in\mathcal{D}(\mathrm{M}_{n})$, the entropy distance $d$ is continuous at $\rho$ if and only if its restriction $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$.
Proof.
Let $\rho\in\mathcal{D}(\mathrm{M}_{n})$. By Lemma 9.6 a), the entropy distance $d$ is continuous at $\rho$ if and only if the rI-projection $\Pi$ is continuous at $\rho$. The inference map
$\Psi^{\mathcal{A}}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathcal{A})\,,\quad\sigma\mapsto\operatorname{argmax}_{\eta\in\pi_{\mathcal{R}}^{-1}(\sigma)}S(\eta)$
has the same values as $\Psi$, whose image $\Psi(\mathcal{D}(\mathcal{R}))$ is included in $\mathcal{D}(\mathcal{A})$ by Ex. 9.3 a). Therefore,
$\Pi=\Psi\circ\pi=\Psi\circ\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$
is continuous at $\rho$ if and only if $\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$, if and only if $\Pi^{\mathcal{A}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$, where
$\Pi^{\mathcal{A}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{A})\,,\quad\Pi^{\mathcal{A}}:=\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}\,.$
Since $\pi_{\mathcal{A}}$ is open by Thm. 1.2, it follows from Lemma 8.1 b) that $\Pi^{\mathcal{A}}\circ\pi_{\mathcal{A}}$ is continuous at $\rho$ if and only if $\Pi^{\mathcal{A}}$ is continuous at $\pi_{\mathcal{A}}(\rho)$. By equation (9.4),
$d|_{\mathcal{D}(\mathcal{A})}(\eta)=S(\Psi^{\mathcal{A}}\circ\pi_{\mathcal{R}}(\eta))-S(\eta)\,,\quad\eta\in\mathcal{D}(\mathcal{A})$
holds. Hence, $\Pi^{\mathcal{A}}$ is continuous at $\pi_{\mathcal{A}}(\rho)$ if and only if $d|_{\mathcal{D}(\mathcal{A})}$ is continuous at $\pi_{\mathcal{A}}(\rho)$, again by Lemma 5.15 1) in [53]. ∎
@@
Prop. 9.8 helps solve the continuity problem of the entropy distance from $\mathcal{E}(\mathcal{R})$ for $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ using the solution of the continuity problem of the restriction $d|_{\mathcal{D}(\mathcal{A})}$ to the real *-subalgebra $\mathcal{A}:=\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R})$ of $\mathrm{M}_{3}$. This solution was obtained from asymptotic curvature estimates [53]. Real *-subalgebras are excluded from Proposition 9.8 but the conclusion is still true, as Ex. 7.2 can replace Thm. 1.2 in the proof of Prop. 9.8.
@@
and continuous at every point in the complement [^6] Unitary similarity with respect to $U:=\exp(\mathrm{i}\frac{\pi}{3\sqrt{3}}(X+Y+Z))$ permutes the Pauli matrices cyclicly. With respect to $U\oplus 1$ it matches $\mathcal{R}$ and the problem considered in [53].. Prop. 9.8 then shows that the entropy distance $d$ is discontinuous at every point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ and continuous at every point in the complement.
This result is consistent with the assertion of Ex. 9.3 c) that the inference map $\Psi$ is discontinuous at $M$ and continuous anywhere else in the ellipse $\mathcal{D}(\mathcal{R})$, as required by Lemma 9.6 b). The points in the set $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{d})$ are explicitly described in Ex. 8.3 b).
Acknowledgments. The author is grateful to Ariel Caticha, Achim Clausing, Chi-Kwong Li, Tim Netzer, Ilya M. Spitkovsky, and Ivan Todorov for remarks, feedback, and discussions. The author gave a lecture on the results of this paper in the special session “Numerical Ranges” of the conference IWOTA 2024 in Canterbury, UK, and wishes to thank the organizers for this opportunity.